OpenStudy (anonymous):

Match each equation with its equivalent equation. Match Term Definition six x equals twenty‾four A) x divided by twelve equals three x divided by six equals six B) four x equals sixteen x divided by two equals sixteen C) x divided by eight equals four eight x equals sixteen D) four x equals eight

OpenStudy (anonymous):

@jigglypuff314

jigglypuff314 (jigglypuff314):

well first you might want to convert it all into math speak >,< :P 6x = 24 A. x/12 = 3 x/6 = 6 B. 4x = 16 x/2 = 16 C. x/8 = 4 8x = 16 D. 4x = 8 so you would solve and match for each...

OpenStudy (anonymous):

yeah but i dont get it what are the answers ?

jigglypuff314 (jigglypuff314):

to solve for x for 6x = 24 divide both sides by 6 so x = 24/6 simplified x/2 = 16 multiply both sides by 2 x = 16*2 x/2 = 16 multiply both sides by 2 8x = 16 divide both sides by 8

OpenStudy (anonymous):

im still not getting them :(

OpenStudy (anonymous):

take for example the first one: 6x = 24 in order to find another one that is the same as this equation what is needed to multiply 6 * ___ = 24

OpenStudy (anonymous):

the answer is 4, so now look at the equations on the right which one gives you a correct value if you put a 4 for x

OpenStudy (anonymous):

that would be 4x = 16 because 4*4 = 16 since the equations match their x value should be the same and will give you a correct answer

OpenStudy (anonymous):

so the first one goes with B

OpenStudy (anonymous):

THE SECOND ONE GOES WITH A?

OpenStudy (anonymous):

the next one is x / 6 = 6, we would need a 36 since 36/6 = 6 now lets use the first one x / 12 = 3 by replacing the x with 36 36/12 = 3 so its true so you answered correct

OpenStudy (anonymous):

THE NEXT ONE GOES WITH d?

OpenStudy (anonymous):

x/2= 16 so you need a 32 so is 32 * 4 =8?

OpenStudy (anonymous):

so it must be the other one lets check 32 / 8 = 4

OpenStudy (anonymous):

YEAH

OpenStudy (anonymous):

3rd is c and 4th is D

OpenStudy (anonymous):

THANK YOU

OpenStudy (anonymous):

np :)